OpenStudy (anonymous):

Solve: 7/8 divided by (-1/4) (show in reduced form)

OpenStudy (anonymous):

-7/2

OpenStudy (anonymous):

7/8 / -1/4 = 7/8 x -4/1 = -28/8 = -7/2